# Fractal geometry predicts dynamic differences in structural and functional connectomes

### v1

@@ 3.2.2 Differences between Structural and Functional connectomes
The first noticeable difference is that, in contrast with equi-M sets for Structural connectomes (which all had a right cusp), all equi-M sets for both Rest and Emotion connectomes have a left cusp. We attribute this to the distribution of negative signs in the Functional connectomes. Further experimental and theoretical work [12] suggests that replacing positive connections with negative connections eventually leads to a change in the orientation of the cusp (process that may involve intermediate stages of the set with multiple cusps). Of course, at what point and in what form this swap occurs is further determined by the placement of the negative connections. Moreover, since the quadratic function is invariant to the sign of the input, adding more negative connections will eventually swap the cusp back to the right. It is very interesting to observe this phenomenon in our data set (especially since it offers great insight into how different dynamics operate on a positive versus signed connectome). Although we do not have a full analytical explanation for it, we have a candidate mechanism that we are investigating in our theoretical work.
@@ 3.2.3 Differences between Rest and Emotion connectomes
First, a basic analysis found direct statistical differences between the graph-theoretical measures for the Rest and Emotion connectomes. This is far from surprising, but is difficult to interpret, since these measures are computed separately for the associated networks. Moreover, differences tended to appear in scattered patterns across the three types of associated networks considered (positive sub-network, negative sub-network and absolute value network). For example, $BC$ was the only measure that was consistently different for the three associated networks (with $p_{pos}<10^{-10}$, $p_{neg}=0.004$ and $p_{abs}<10^{-10}$, respectively). In addition to this, the positive sub-network exhibited significant differences in motive strengths $M1$ and $M2$ (with $p<10^{-10}$ and $p=0.059$, respectively), the absolute value network shows trend differences in $GE$ and $TR$ ($p=0.059$ and $p=0.045$), and the negative network captured differences in all measures except $A$ and $D$ (with $p<0.05$ across the board). We believe these patterns may be important, but it is very hard to judge their impact, since none of them encompasses the interconnections between the positive and negative sub-networks in assembling the whole connectome.
One can in fact use the equi-M set to further assess whether/to what extent any of the three associated networks capture the dynamic flow in the original connectome. To do so, we computed the equi-M set for each of the 48 connectomes, for all three associated networks. In all cases, the resulting equi-M sets were all just scaled copies of the traditional equi-M set (the sets for the absolute networks are shown in Figures 14 and 16, for illustration). This is in itself a powerful result. First, it suggests that the dynamics propagates completely differently in the associated networks when considered separately than in the original network. Second, it shows that significant differences in the structure of these associated networks may lead to indistinguishable dynamics when considered independently. Moreover, the striking differences between the equi-M sets for the original connectomes were completely wiped out and were devoid of network-derived information (all were scaled Mandelbrot sets). Our analysis of the Structural data ensures that this is not the case for all positive connectomes, since the Structural connectomes themselves produced a wide diversity of equi-M sets. The loss of dynamic diversity must therefore be due to the loss of information on a signed network’s architecture when considering only parts of it, or when ignoring the signs. Third, this tells us that the statistical differences found between the associated networks for Rest and Emotion connectomes should not be used to assess the impact of the whole connectome on dynamics.
@@
Figure 7: Different approaches to the Rest connectome data using the equi-M set. (a) Statistical equi-M set (the light to dark gradient shows the fraction of equi-M sets that the respective point in $\mathbb{C}$ belongs to), together with the collection of all equi-M sets derived from Rest connectomes (overlaid in gray). (b) The equi-M set computed for the prototypical Rest connectome.
Figure 8: Different approaches to the Emotion connectome data using the equi-M set. (a) Statistical equi-M set (the light to dark gradient shows the fraction of equi-M sets that the respective point in $\mathbb{C}$ belongs to), together with the collection of all equi-M sets derived from Emotion connectomes (overlaid in gray). (b) The equi-M set computed for the prototypical Emotion connectome.
For a visual assessment of each individual group,  Figures 7 and 8 illustrate the stochastic and the prototypical equi-M set for Rest and Emotion connectomes respectively. One may notice that both statistical and prototypical sets for Emotion are a lot more spread out than those for Rest. This is interesting since it suggests more diverse functional dynamics when processing emotional content and more unifying dynamics when processing neutral input (possibly related to reliance on the default network at rest). To verify this interesting result quantitatively, we sampled the spread of the boundaries of the equi-M set in each case, by taking a rotating radius and keeping track of where it intersects these boundaries. We obtained a distribution $D_{\theta k}$, where $0\leq\theta\leq 2\pi$ is the sampling angle, and $1\leq k\leq 48$ is the subject in the data set. In Figure 9 we compare the Rest and Emotion histograms for $D_{\theta k}$ (considered for all values of $\theta$ and $k$), as well as for the means $\langle D_{\theta k}\rangle$ of these distributions over all sampled $\theta$s. A Levine non-parametric test shows that, while the means of these distributions are not a distinguishing factor between them, they are different in variance. For the overall distributions in the left figure panel, we obtained significantly different variances $\sigma_{\text{Rest}}=0.0035$ and $\sigma_{Emotion}=0.0044$, $F=1700$, with $df_{Rest}=1$, $df_{Emotion}=96094$ and $p<10^{-4}$. For the mean distribution in the right panel, averaging somewhat obscured the effect, but we still found a trend in differences, with $\sigma_{\text{Rest}}=0.0028$ and $\sigma_{Emotion}=0.0038$, $F=3.02$, $df_{Rest}=1$, $df_{Emotion}=94$ and $p=0.08$). This confirms a generally wider variability on dynamics for Emotional task than for Rest.
@@
A natural next step is to revisit more systematically the relationship between the measures of dynamics (geometric and Fourier measures) and the graph-theoretical measures (for the connectomes’ associated networks). By comparing the correlation profiles in Figure 5a with those in Figure 10, one can easily see some similarities, but also crucial differences.
Figure 10: Correlations between the graph-theoretical measures of the associated networks, and the geometric measures of the equi-M set of the original network. The left panels represent Rest connectomes, and the right panels represent Emotion connectomes. The correlations are presented, top to bottom, for: positive subnetwork, negative subnetwork and absolute value network. In each panel, the measures are listed along both coordinate axes, in the order found in the tables, and in the order introduced in the text. For $C$, $(R)$ and $G$, we used their absolute values, to illustrate the distance to the origin. For the complex modes $m_{j}$, with $-2\leq j\leq 2$, we use their moduli (to illustrate their strength). In this scheme, the three diagonal squares show correlations between measures within the same category, and the off-diagonal rectangles show correlations between measures in different categories. For simplicity, we only included the $\rho$ values here.
@@
Second, the profile of correlations between the geometric measures and the graph-theoretical measures for the associated networks is lower. In the case of Rest connectomes, significant correlations persist primarily between the subsets of measures that are correlated with each other within their own group. The top left panel in Figure 10 shows that, for the positive sub-network, there is a subset of graph measures and a subset of geometric measures that are both “unified” by high positive mutual correlations within each set (red blocks). In addition, these are the measures that also show significant negative inter-correlations (blue stripes). The same is true, to a lesser extent, of the negative sub-network and the absolute value network for Rest connectomes. This suggests that some measures of the “size” of the equi-M set may still capture graph-theoretical graph patterns in the associate networks of the connectome. However, our experiment of constructing equi-M sets independently for each associate network led to equi-M sets which were identical up to scaling to the traditional Mandelbrot set, across the whole data. This further tells us that the only relevance that the architecture of these networks may have on the global network dynamics is weak (it only impacts the size of the equi-M set, and none of the more subtle shape aspects). This idea is reinforced by the fact that the only significant correlations for Fourier-derived measures appear for the perimeter, which is also a measure of size, and not the modes (which are finer descriptions of the boundary variability of the equi-M set). These effects are even more pronounced in the case of Emotion connectomes (right panels).
@@
To illustrate this, the panels of Figure 11 offer two perspectives on the differences between the modes for Structural vs Functional connectomes, and also between the modes for Emotion vs Rest connectomes. The first panel shows the log-log power spectra for all connectomes, as well as the group averages. Notice that the magnitudes of the modes for the Structural data are dramatically larger than those for the Functional data (both Rest and Emotion). This is confirmed by the second panel, which illustrates as a specific example the distribution of the second mode $m_{2}$ in the complex plane, as separate scatter plots accompanied by the corresponding confidence ellipses. The distribution for the Structural data (orange dots) is dramatically more spread out than the distributions for Rest (blue) and Emotion (pink dots), with Levine test significances $df<10^{-4}$ for both Structural vs. Rest and Structural vs. Emotion.
Notice that the differences between modes for Rest and Emotion connectomes are not as dramatic, but still powerful. Although the Rest vs. Emotion power spectra do not appear significantly different, the distributions of each of these modes in $\mathbb{C}$ present quantifiable differences between Rest and Emotion. These differences are illustrated for $m_{2}$ and $m_{3}$ in particular in the bottom tow panels. In each case, confidence ellipses are included, emphasizing the differences between their spread and principal components. In addition, a separate Levene test confirms that the distribution of $m_{2}$ for Rest have significantly lower variability ($\sigma_{Rest}=0.05$) than that for Emotion ($\sigma_{Emotion}=0.11$), with $F=6.70$ and $p=0.01$. Similarly for $m_{3}$, $\sigma_{Rest}=0.02$ and $\sigma_{Emotion}=0.05$, with $F=8.78$ and $p=0.003$. This confirms the earlier result that the shape variability of the equi-M sets is higher in the Emotion task than in Rest.
@@ 4 Discussion
This was done in two ways. First, in addition to computing the equi-M sets for whole Functional connectomes, we also computed the sets corresponding to the associated networks. Although the original equi-M sets exhibited a lot of variability in shape (and not just size), the sets for the associated networks were identical to the Mandelbrot set up scaling (via small detail potentially due to precision). This made it clear that, for the quadratic model, many of the subtleties in dynamics that encode differences between individuals cannot be captured simply by considering the associated networks. It is not too big a stretch to suggest that this may also be the case for more complex node dynamics (such as neural dynamics). This effect did not come from simply transitioning from a signed Functional connectome to a positive network (since Structural connectomes, although positive, also have shape diversity). The effect is likely due to the sparse distribution of connections comprised in the positive and negative sub-networks, which, taken independently, are insufficient to recover the complexity of the overall connectome, in absence of the interactions between them.
@@
Indeed, our final analysis showed Fourier-based modes to be an efficient differentiator not only between Structural and Functional equi-M sets, but also between our two types of Functional connectomes, in that the modes of the Emotion connectomes were significantly more spread out than those of Rest connectomes. This, together with the equi-M sets themselves showing a higher spread in Emotion than in Rest, suggested increased variability and diversity in dynamics for Emotional task than for Rest. Aside form its relevance as a quantitative assessment, this is an interesting result in the neuroscience context. It is not surprising that subjects appear to have more variable dynamics in Emotional processing than in Resting state (since Rest is known to rely primarily on the default network, while Emotion tasks may engage a variety of mechanisms, different for each subject). What is surprising and heartening is that the Fourier measures of the equi-M set were able to capture this idea.
@@ Acknowledgements
The project received support from an AMS-Simons PUI Faculty Grant (Rădulescu) and from the UEFISCDI, project number ROSUA-2024-0002 (Rădulescu, Kaslik, Fikl). Michael Anderson’s work on the project was supported by New Paltz RSCA Summer and Academic Year Experience for Undergraduate Programs.
Appendix A Equi-M sets for all subjects
Figure 12: Structural equi-M sets for all subjects in our data set. All sets are represented in the complex region $[-1,0.3]\times[-0.65,0.65]$. The equi-M set in each panel appears in black, and the colors represent the number of iterations needed by each point in order to exit the escape disc.
Figure 13: Equi-M sets for Rest functional connectomes. The sets present with a wide variety of shapes across the data set.
Figure 14: Equi-M sets for Rest functional connectomes without signs. When using the absolute value of the functional connectome as a generating network, the resulting sets are scaled copies of the traditional Mandelbrot set.
Figure 15: Equi-M sets for Emotion functional connectomes. The sets show a wide variety of shapes across the data set.
Figure 16: Equi-M sets for Emotion functional connectomes without signs. When using the absolute value of the functional connectome as a generating network, the resulting sets are again just scaled copies of the traditional Mandelbrot set.

### v2

@@ 3.2.2 Differences between Structural and Functional connectomes
We first examine differences in the equi-M sets between structural and functional connectomes, as seen in Figures 1213 and 15 in the Appendices. The first noticeable difference is that, in contrast with equi-M sets for Structural connectomes (which all had a right cusp), all equi-M sets for both Rest and Emotion connectomes have a left cusp. We attribute this to the distribution of negative signs in the Functional connectomes. Further experimental and theoretical work not included in this study [12] suggests that replacing positive connections with negative connections eventually leads to a change in the orientation of the cusp (process that may involve intermediate stages of the set with multiple cusps). Of course, at what point and in what form this swap occurs is further determined by the placement of the negative connections. Moreover, since the quadratic function is invariant to the sign of the input, adding more negative connections will eventually swap the cusp back to the right. It is very interesting to observe the left-oriented cusps for the Functional connectomes in our data set (especially because it offers great insight into how different dynamics operate on a positive versus signed connectome). Although we do not have a full analytical explanation for it, we have a candidate mechanism that we are investigating in our theoretical work.
@@ 3.2.3 Differences between Rest and Emotion connectomes
First, a basic analysis (using the Wilcoxon rank sum test) found direct statistical differences between the graph-theoretical measures for the Rest and Emotion connectomes. This is far from surprising, but is difficult to interpret, since these measures are computed separately for the associated networks. Moreover, differences tended to appear in scattered patterns across the three types of associated networks considered (positive sub-network, negative sub-network and absolute value network). For example, $BC$ was the only measure that was consistently different for the three associated networks (with $p_{pos}<10^{-10}$, $p_{neg}=0.004$ and $p_{abs}<10^{-10}$, respectively). In addition to this, the positive sub-network exhibited significant differences in motif strengths $M1$ and $M2$ (with $p<10^{-10}$ and $p=0.059$, respectively), the absolute value network shows trend differences in $GE$ and TRTR ($p=0.059$ and $p=0.045$), and the negative network captured differences in all measures except $A$ and $D$ (with $p<0.05$ across the board). We believe these patterns may be important, but it is very hard to judge their impact, since none of them encompasses the interconnections between the positive and negative sub-networks in assembling the whole connectome.
One can in fact use the equi-M set to further assess whether/to what extent any of the three associated networks capture the dynamic flow in the original connectome. To do so, we computed the equi-M set for each of the 48 connectomes, for all three associated networks. In all cases, the resulting equi-M sets were all just scaled copies of the traditional equi-M set (the sets for the absolute networks are shown in Figures˜14 and 16, for illustration). This is in itself a powerful result. First, it suggests that the dynamics propagates completely differently in the associated networks when considered separately than in the original network. Second, it shows that significant differences in the structure of these associated networks may lead to indistinguishable dynamics when considered independently. Moreover, the striking differences between the equi-M sets for the original connectomes were completely wiped out and were devoid of network-derived information (all were scaled Mandelbrot sets). Our analysis of the Structural data ensures that this is not the case for all positive connectomes, since the Structural connectomes themselves produced a wide diversity of equi-M sets. The loss of dynamic diversity must therefore be due to the loss of information on a signed network’s architecture when considering only parts of it, or when ignoring the signs. Third, this tells us that the statistical differences found between the associated networks for Rest and Emotion connectomes should not be used to assess the impact of the whole connectome on dynamics.
@@
Figure 7: Different approaches to the Rest connectome data using the equi-M set. (a) Statistical equi-M set: the light to dark gradient shows the fraction of equi-M sets that the respective point in $\mathbb{C}$ belongs to. (b) The equi-M set computed for the prototypical Rest connectome.
Figure 8: Different approaches to the Emotion connectome data using the equi-M set. (a) Statistical equi-M set: the light to dark gradient shows the fraction of equi-M sets that the respective point in $\mathbb{C}$ belongs to. (b) The equi-M set computed for the prototypical Emotion connectome.
[…]
@@
A natural next step is to revisit more systematically the relationship between the measures of dynamics (geometric and Fourier measures) and the graph-theoretical measures (for the connectomes’ associated networks). By comparing the correlation profiles in Figure˜5a with those in Figure˜10, one can easily see some similarities, but also crucial differences.
Figure 10: Correlations between the graph-theoretical measures of the associated networks, and the geometric measures of the equi-M set of the original network. The left panels (a,c,e) represent Rest connectomes, and the right panels (b,d,f) represent Emotion connectomes. The correlations are presented, top to bottom, for: positive subnetwork, negative subnetwork and absolute value network. In each panel, the measures are listed along both coordinate axes, in the order found in the tables, and in the order introduced in the text. For $C$, Re(R)(R) and $G$, we used their absolute values, to illustrate the distance to the origin. For the complex modes $m_{j}$, with $2\leq j\leq 2$, we use their moduli (to illustrate their strength). In this scheme, the three diagonal squares show correlations between measures within the same category, and the off-diagonal rectangles show correlations between measures in different categories. For simplicity, we only included the $\rho$ values here.
@@
Second, the profile of correlations between the geometric measures and the graph-theoretical measures for the associated networks is lower. In the case of Rest connectomes, significant correlations persist primarily between the subsets of measures that are correlated with each other within their own group. The top left panel in Figure˜10 shows that, for the positive sub-network, there is a subset of graph measures and a subset of geometric measures that are both “unified” by high positive mutual correlations within each set (red blocks). In addition, these are the measures that also show significant negative inter-correlations (blue stripes). The same is true, to a lesser extent, of the negative sub-network and the absolute value network for Rest connectomes. This suggests that some measures of the “size” of the equi-M set may still capture graph-theoretical graph patterns in the associate networks of the connectome. However, our experiment of constructing equi-M sets independently for each associate network led to equi-M sets which were identical up to scaling to the traditional Mandelbrot set, across the whole data. This further tells us that the only relevance that the architecture of these networks may have on the global network dynamics is weak (it only impacts the size of the equi-M set, and none of the more subtle shape aspects). This idea is reinforced by the fact that the only significant correlations for Fourier-derived measures appear for the perimeter, which is also a measure of size, and not the modes (which are finer descriptions of the boundary variability of the equi-M set). These effects are even more pronounced in the case of Emotion connectomes (right panels).
@@
To illustrate this, the panels of Figure˜11 offer two perspectives on the differences between the modes for Structural vs Functional connectomes, and also between the modes for Emotion vs Rest connectomes. The first panel shows the log-log power spectra for all connectomes, as well as the group averages. Notice that the magnitudes of the modes for the Structural data are dramatically larger than those for the Functional data (both Rest and Emotion). This is confirmed by the second panel, which illustrates as a specific example the distribution of the second mode $m_{2}$ in the complex plane, as separate scatter plots accompanied by the corresponding confidence ellipses. The distribution for the Structural data (orange dots) is dramatically more spread out than the distributions for Rest (blue) and Emotion (pink dots), with Levene test significances $p<10^{-4}$ for both Structural vs. Rest and Structural vs. Emotion.
Notice that the differences between modes for Rest and Emotion connectomes are not as dramatic, but still powerful. Although the Rest vs. Emotion power spectra do not appear significantly different, the distributions of each of these modes in $\mathbb{C}$ present quantifiable differences between Rest and Emotion. These differences are illustrated for $m_{2}$ and $m_{3}$ in particular in the bottom two panels. In each case, confidence ellipses are included, emphasizing the differences between their spread and principal components. In addition, a separate Levene test confirms that the distribution of $m_{2}$ for Rest have significantly lower variability ($\sigma_{Rest}=0.05$) than that for Emotion ($\sigma_{Emotion}=0.11$), with $F=6.70$ and $p=0.01$. Similarly for $m_{3}$, $\sigma_{Rest}=0.02$ and $\sigma_{Emotion}=0.05$, with $F=8.78$ and $p=0.003$. These results are significant even when using a Bonferroni corrected $\alpha=0.01$ to account for using comparisons over five modes. This confirms the earlier result that the shape variability of the equi-M sets is higher in the Emotion task than in Rest.
@@ 4 Discussion
This was done in two ways. First, in addition to computing the equi-M sets for whole Functional connectomes, we also computed the sets corresponding to the associated networks. Although the original equi-M sets exhibited a lot of variability in shape (and not just size), the sets for the associated networks were identical to the Mandelbrot set up scaling (via small detail potentially due to precision). This made it clear that, for the quadratic model, many of the subtleties in dynamics that encode differences between individuals cannot be captured simply by considering the associated networks.
It is not too big a stretch to suggest that this may also be the case for more complex node dynamics (such as neural dynamics). This effect did not come from simply transitioning from a signed Functional connectome to a positive network (since Structural connectomes, although positive, also have shape diversity). The effect is likely due to the sparse distribution of connections comprised in the positive and negative sub-networks, which, taken independently, are insufficient to recover the complexity of the overall connectome, in absence of the interactions between them. This hints towards the fact that the interaction between the positively and negatively correlated regions contribute fundamentally to the emergent properties of the network, rather than simply being an artifact of its structural composition.
@@
Indeed, our final analysis showed Fourier-based modes to be an efficient differentiator not only between Structural and Functional equi-M sets, but also between our two types of Functional connectomes, in that the modes of the Emotion connectomes were significantly more spread out than those of Rest connectomes. This, together with the equi-M sets themselves showing a higher spread in Emotion than in Rest, suggested increased variability and diversity in dynamics for Emotional task than for Rest. Aside form its relevance as a quantitative assessment, this is an interesting result in the neuroscience context. It is not surprising that subjects appear to have more variable dynamics in Emotional processing than in Resting state (since Rest is known to rely primarily on the default network [39, 40, 41], while Emotion tasks may engage a variety of mechanisms, different for each subject). What is surprising and heartening is that the Fourier measures of the equi-M set were able to capture this idea.
@@ Acknowledgements
The project received support from an AMS-Simons PUI Faculty Grant (Rădulescu) and from a joint NSF-UEFISCDI grant: Project No. #2408407 (Rădulescu), Project No. ROSUA-2024-0002 (Kaslik, Fikl). Michael Anderson’s work on the project was supported by New Paltz RSCA Summer and Academic Year Experience for Undergraduate Programs.
Data Availability
The software developed and used in this study is openly available in the Zenodo repository at https://doi.org/10.5281/zenodo.16921908 under the MIT license. This archive (version 1.1) includes all source code and documentation required to generate the equi-M sets.
Appendix A: Equi-M sets for all subjects
Figure 12: Structural equi-M sets for all subjects in our data set. All sets are represented in the complex region $1,0.3]\times[-0.65,0.65]$, in resolution $1200\times 1200$. The equi-M set in each panel appears in black, and the colors represent the number of iterations needed by each point in order to exit the escape disc.
Figure 13: Equi-M sets for Functional Rest connectomes for all subjects in our data set. All sets are represented in the complex region $0.03,0.03]\times[-0.03,0.03]$, in resolution $1200\times 1200$. The equi-M set in each panel appears in black, and the colors represent the number of iterations needed by each point in order to exit the escape disc. The sets show a wide variety of shapes across the data set.
Figure 14: Equi-M sets for the absolute value associated networks of the Functional Rest connectomes. All sets are represented in the complex region $0.002,0.0005]\times[-0.0015,0.0015]$, in resolution $1200\times 1200$. The equi-M set in each panel appears in black, and the colors represent the number of iterations needed by each point in order to exit the escape disc. When using the absolute value of the functional connectome as a generating network, the resulting sets are scaled copies of the traditional Mandelbrot set.
Figure 15: Equi-M sets for Functional Emotion connectomes for all subjects in our data set. All sets are represented in the complex region $0.03,0.03]\times[-0.03,0.03]$, in resolution $1200\times 1200$. The equi-M set in each panel appears in black, and the colors represent the number of iterations needed by each point in order to exit the escape disc. The sets show a wide variety of shapes across the data set.
Figure 16: Equi-M sets for the absolute value associated networks of the Functional Emotion connectomes. All sets are represented in the complex region $0.002,0.0005]\times[-0.0015,0.0015]$, in resolution $1200\times 1200$. The equi-M set in each panel appears in black, and the colors represent the number of iterations needed by each point in order to exit the escape disc. When using the absolute value of the functional connectome as a generating network, the resulting sets are scaled copies of the traditional Mandelbrot set.